# Training on the Test Task Confounds Evaluation and Emergence

### v1

@@ 3.1 Fine-tuning on the test task
For this section, we only consider models trained before November 2023, since we hypothesize that older models do not train on the test task much. We randomly split models into two cohorts: a control group and a treatment group. We fine-tune the treatment group on the datasets of task-relevant data introduced in Section 2. We fine-tune on each dataset independently, for a single epoch. We then evaluate the benchmark performance of the two cohorts, as well as their performance after adjusting for test task training. As in the previous section, we adjust for test task training by fine-tuning all models on the test task before evaluation.
We plot in Figure 3 the two cohorts’ benchmark performance before and after the adjustment. We repeat the statistical analysis of Section 2.2 and report the estimated coefficient indicating the average difference in benchmark performance between the two cohorts when controlling for compute.
Fine-tuning the treatment group results in large differences in performance between the control group and the treatment group, see Figure 3 middle. Qualitatively, the differences between the control and treatment group resemble those observed between newer and older models in Section 2.2. In particular, the fine-tuned models Pareto dominate the non fine-tuned models. Quantitatively, the estimated increase in performance due to fine-tuning is statistically significant and large. Importantly, it is also similar to the difference in performance between newer and older models estimated in Section 2.2. Therefore, fine-tuning older models on the test task gives rise to qualitatively and quantitatively similar confounding to that observed between newer and older models. This is consistent with our running hypothesis that newer models are largely equivalent to older models that trained on the test task.
After adjusting for test task training by further fine-tuning both the control and treatment groups on the test task, we observe that models in the treatment group are no longer outliers in terms of performance-per-compute, see Figure 3 right. Quantitatively, the estimated increase in performance is both small and not statistically significant. We therefore validate a vital soundness property of the proposed adjustment procedure: after deliberately training some models on the test task, we can undo their advantage over other models by further training all models on the test task.
@@ 3.2 Reformulating the test task
In this section we consider two additional benchmarks from the HF leaderboard: ARC Challenge (Clark et al., 2018) and HellaSwag (Zellers et al., 2019). Similarly to MMLU, ARC is comprised of grade-school level questions. HellaSwag instead tests for commonsense reasoning. Like MMLU, the questions in ARC and HellaSwag are accompanied by four possible answers, among which the model must differentiate the correct one. The standard MMLU evaluation formulates questions as multiple-choice: all four answer choices are listed, and the model is promoted to pick one. In contrast, ARC and HellaSwag use “cloze” evaluations: a models’ answer is taken to be that with the largest completion likelihood given the input question.
We evaluate all models on ARC and HellaSwag using the standard cloze evaluation, and plot their benchmark performance in Figure 4 left. We repeat the statistical analysis of Section 2.2, and report the average difference in performance between newer and older models after controlling for pretraining compute. Qualitatively, we observe that older models and newer models have very similar scaling trends. Quantitatively, the estimated difference in performance between newer and older models is small and not statistically significant. That is, newer models do not outperform older models on ARC and HellaSwag.
We then reformulate ARC and HellaSwag as MMLU-style multiple-choice questions, and plot the resulting benchmark performance in Figure 4 center. We observe large differences in performance between newer and older models. Qualitatively, these differences in performance resemble those observed for MMLU. In particular, newer models Pareto dominate in terms of performance-per-compute. Quantitatively, we find the difference in performance between newer and older models to be significant, positive, and large, and to be roughly similar in magnitude to that estimated for MMLU in Section 2.2. Therefore, reformulating the test task as multiple choice question answering leads to similar confounding to that observed for MMLU. This suggest that what causes the outliers in MMLU is likely not memorization of specific testing data (i.e., due to data contamination or leakage), but rather an improved ability for MMLU-style prompts.
We adjust for test task training by fine-tuning all models on the MMLU auxiliary training set, and plot their ARC Challenge and HellaSwag scores in Figure 4. We observe that newer models are no longer outliers in terms of performance-per-compute. Moreover, we no longer find evidence of a significant difference in performance between newer and older models. The proposed adjustment is therefore effective in removing the confounding resulting from newer models overperforming for MMLU-style prompts.
@@
We evaluate MMLU using the “cloze” methodology instead of the usual multiple-choice prompts. We plot the results in Figure 5 center. With cloze evaluations, newer models are no longer are outliers in terms of MMLU performance. In fact, the difference in performance between newer and older models is now small and not statistically significant. This suggests that the standard MMLU evaluation conflates knowledge-testing with testing a models’ ability to answer multiple choice questions. For instance, smaller non fine-tuned models suffer from particularly strong A-bias for multiple-choice questions (Dominguez-Olmedo et al., 2023). Newer models therefore attain higher MMLU scores largely because they are better at multiple-choice question answering, and not because they necessarily “know more”.
@@ 4 Implications for model comparisons
Our findings indicate that training on the test task acts as a major confounder of LLM benchmark evaluations. We now discuss its implications for the relative comparison of model families (Section 4.1), as well as its implications for measuring progress in model capabilities over time (Section 4.2).
@@ 4.1 Comparing model families
We compare the MMLU and GSM8K performance of the Pythia, Llama 2, and Qwen 1.5 model families, which likely train on the test task to very different extents. Pythia was trained on the Pile (Gao et al., 2020), a collection of curated datasets that are unlikely to contain much test task data. Llama 2 was trained mostly on web data, which is reasonable to assume may contain test task data. Lastly, Qwen 1.5 explicitly includes instruction data in its pretraining mixture, thus likely training on the test task to a large extent.
In Figure 6 we plot the MMLU and GSM8K scores of the Llama 2, Qwen 1.5, and Pythia families of models, as well as their adjusted accuracy (i.e., after fine-tuning on task relevant data). Without adjustment, Qwen 1.5 appears to be the superior model family: it Pareto dominates both the Llama 2 and Pythia models. Furthermore, all Pythia models perform no better than random chance, and thus it is unclear what benefit scaling Pythia might bring. After fine-tuning the models on the the test task, however, all three model families exhibit very similar scaling trends. Therefore, when correcting for the confounding introduced by test task training it is unclear if any of the model families is superior to the others beyond their pretraining compute.
Interestingly, recent work equates pretraining data quality with downstream benchmark performance (Penedo et al., 2024; Li et al., 2024). For example, the Pile (Pythia’s pretraining dataset) is thought to be inferior to filtered web data. In light of our findings, it is plausible that a major contributing factor to the superior performance of “higher quality” pretraining datasets is that they contain a larger share of test task data.
@@ 4.2 Progress in model capabilities
One purpose of benchmarks is to track progress in model capabilities. In Figure 7 we plot the Pareto frontier of benchmark accuracy against pretraining compute, both for models trained before November 2023 and for all models. We measure progress by considering the area of improvement of the Pareto frontier since November 2023, shaded in green. Without adjustment, the difference between the two Pareto frontiers is rather large for both MMLU and GSM8K, indicating substantial progress since November 2023. After fine-tuning models on the test ask, however, the area of improvement reduces by a sixfold. Therefore, the confounding introduced by test task training leads to substantially overestimating the progress in MMLU and GSM8K capabilities per unit of compute achieved by recent model families.
On the other hand, recent models tend to be trained on more data. Given the Chinchilla scaling laws (Hoffmann et al., 2022), it is remarkable that newer, smaller models match the performance of older, larger ones for the same amount of pretraining compute. Since inference and fine-tuning of smaller models is substantially cheaper, recent models can be much more accessible to less well-resourced institutions, with little cost in performance. For example, we find that Llama 3 8B closely matches the performance of Llama 2 70B.
@@ 5 Implications for emergence
Throughout our evaluations, we observe emergent behaviour for MMLU and GSM8K: models perform at near random chance up to a certain scale of pretraining compute, followed by relatively sharper improvements in performance at larger scales (Wei et al., 2022). After training on the test task, however, emergence for MMLU and GSM8K appears to occur at substantially lower scales. We dedicate this section to more closely investigate the relationship between training on the test task and emergence.
@@
We consider only models trained before November 2023, as we have established that these models train on the test task to a lesser extent. We evaluate the models at intermediate checkpoints as they train on the datasets of task relevant data introduced in Section 2.1. We fit and in Equation 1 to the different intermediate checkpoints, and report in Figure 8 the corresponding points of emergence . We find that emergence arises at increasingly lower compute regimes as models increasingly train on the test task. For instance, for MMLU the non-fine tuned models exhibit emergence at around FLOPs, roughly the scale of Pythia 6.9B. After training on 64,000 examples, emergence arises around around FLOPs, that is, roughly the scale of Pythia 410M. That is, the benchmark performance of models after training on the test task is predictable at substantially lower scales.
@@
The log-linear relationship between pretraining loss and compute is well-established (Kaplan et al., 2020; Hoffmann et al., 2022). We observe that, for the compute ranges that we consider, training on the test task increasingly recovers log-linear scaling between pretraining compute and benchmark accuracy. Similarly to the earlier section, we evaluate intermediate checkpoints but instead fit log-linear functions in Figure 9. We observe that the R2 of the fit improves substantially as the models train on more task-relevant data. For MMLU, the value jumps from to after training on 64,000 examples. Therefore, after training on the test task almost all of the variation in accuracy can be explained by log-linear scaling of pre-training compute.
Discussion.
Schaeffer et al. (2024a) argue that emergence appears due to the choice of metric. To mitigate emergence, they suggest to consider Brier score instead of accuracy. We observe, however, that the emergent behaviour of MMLU does not disappear when using the Brier score, see Figure 5 right, nor that of ARC and HellaSwag when framed as multiple-choice questions, see Figure 16 in Appendix D. While more complex changes of metric might resolve the emergence in multiple-choice QA (Schaeffer et al., 2024b), we discuss two practical solutions to obtain predictive scaling while maintaining accuracy as the evaluation metric.
For MMLU and multiple-choice benchmarks more broadly, we consistently observe that cloze evaluations yield smoother and more predictable scaling even when using accuracy as the evaluation metric. Since the purpose of these benchmarks is knowledge-testing more so than testing multiple-choice answering ability, cloze evaluations should be preferable insofar predictive scaling is an important consideration.
More broadly, if sufficient task relevant data is available, then training on the test task can result in much more predictable scaling by shifting emergence to smaller compute scales. Crucially, the evaluation metric and methodology need not be changed. Note that in many settings it is not a priori apparent what metric or evaluation methodology might results in predictive scaling. Scaling laws after fine-tuning correspond to those of more “specialist” models, which for some domains –such as the legal domain (Dominguez-Olmedo et al., 2024)– or purposes –e.g., safety– might be preferable to the scaling law of generalist models.
@@ 6 Related work
Data contamination or test-set contamination refers to any overlap between the training and the test data such that test results overestimate a model’s generalization performance. The scale and often little curation of present-day pretraining corpora exacerbates data contamination concerns in language model evaluations (Jiang et al., 2024). Consequently, data contamination is usually discussed in the technical reports accompanying model releases (Radford et al., 2019; Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b). However, detecting and preventing data contamination is currently an open problem (Gunasekar et al., 2023; Yang et al., 2023b; Golchin and Surdeanu, 2023). Roberts et al. (2023) and Li and Flanigan (2024) find that models often perform better on datasets that were publicly available during model training. While almost all models that we consider were released after MMLU and GSM8K, we nonetheless find that, controlling for compute, more recent models perform better. These performance gains are unlikely to be driven solely by test set leakage and require additional explanation.
@@
The effectiveness of fine-tuning on the training set accompanying LLM benchmarks is well-known (Wei et al., 2021; Wang et al., 2022; Chung et al., 2024). Consequently, many influential instruction-tuning datasets contain or are partly derived from benchmark train data (Wei et al., 2021; Honovich et al., 2022; Mukherjee et al., 2023). Li and Flanigan (2024) identify small amounts of benchmark-specific data in the publicly available Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) instruction-tuning sets. Zhou et al. (2023b) empirically analyze the effects of fine-tuning on benchmark-specific data and warn about its impacts on benchmark validity. To circumvent these issues, recent work has focused on indirect indicators of broader data contamination, such as a lack of robustness to task transformations (Wu et al., 2023), or underperformance on benchmarks with novel task combinations (Yu et al., 2023a). In contrast, we find evidence for training on the test task without the need for explicitly identifying specific data points used at training time, or modifying tasks. In addition, our method allows us to quantify and correct for the effects of training on the test task on benchmark performance.
@@
Emergent capabilities (Wei et al., 2022; Ganguli et al., 2022) refer to levels of model performance at large scales that cannot be easily predicted by extrapolating from smaller scales. Wei et al. (2022) report emergent capabilities for various benchmarks including MMLU and GSM8K (Srivastava et al., 2022). However, Srivastava et al. (2022); Schaeffer et al. (2024b) find that the log-probability of the correct answer often improves smoothly, even when other metrics seem to show emergence. Lu et al. (2023) argue that most emergent capabilities can be explained by in-context-learning. Schaeffer et al. (2024a) argue that emergent capabilities are mostly an artifact of non-linear and discontinuous evaluation metrics like accuracy. In contrast, we find signs of emergence on tasks like MMLU, even when using continuous metrics like the Brier score. We additionally show that fine-tuning on the test task yields more predictive scaling by shifting the point of emergence to substantially smaller compute scales.

### v2

@@ 3.1 Fine-tuning on the test task
For this section, we only consider models trained before November 2023. We split models into two cohorts: a control group and a treatment group. We take these older models as a control group. We then create a treatment group by fine-tuning the control group on the datasets of task-relevant data introduced in Section 2. We only fine-tune models with at least FLOPs, the pre-training compute of the smallest newer model. We fine-tune for a single epoch. We plot in Figure 3 top the benchmark performance of the two cohorts.
Qualitatively, the differences in performance between the control and treatment groups resembles the differences observed between newer and older models, contrast Figure 3 with Figure 1. Quantitatively, the estimated performance gain from fine-tuning is similar to the difference between newer and older models estimated in Section 2.2. That is, fine-tuning older models on the test task produces both qualitatively and quantitatively similar confounding to that observed between newer and older models. This results further supports our running hypothesis that newer models are largely equivalent to older models that have trained on the test task. They also demonstrate the large effect that training on the test task can have on benchmark performance. Note that the gain in performance of the treatment group is slightly larger than the difference in performance between newer and older models. This is to be expected, as all models in the treatment group are fine-tuned on the test task, whereas not all new models may train on the test task.
We then apply our proposed adjustment by further fine-tuning both the control and treatment groups on the test task, see Figure 3 right. After the adjustment, the estimated difference in performance between the control and treatment group is both small and not statistically significant. We therefore validate a vital soundness property of the proposed adjustment procedure: after deliberately training some models on the test task, we can undo their advantage over other models by further training all models on the test task.
@@ 3.2 Reformulating the test task
In this section we show that reformulating other benchmarks as multiple-choice question answering tasks leads to similar differences in performance between older and newer models. We consider two additional benchmarks from the HF leaderboard v1: ARC Challenge (Clark et al., 2018) and HellaSwag (Zellers et al., 2019). Similarly to MMLU, ARC is comprised of grade-school level questions. HellaSwag instead tests for commonsense reasoning. Like MMLU, the questions in ARC and HellaSwag are accompanied by four possible answers, among which the model must differentiate the correct one. ARC and HellaSwag use “cloze” evaluations: a models’ answer is taken to be that with the largest completion likelihood given the input question. In contrast, MMLU formulates questions as multiple-choice: all four answer choices are listed, and the model is promoted to pick one of the answer choices.
We first evaluate all models on ARC and HellaSwag using the standard cloze evaluation, and plot their benchmark performance in Figure 4 left. We repeat the statistical analysis of Section 2.2. We find that the estimated difference in performance between newer and older models is small and not statistically significant. That is, newer models do not outperform older models on ARC and HellaSwag.
We then reformulate ARC and HellaSwag as MMLU-style multiple-choice questions, and plot the resulting benchmark performance in Figure 4 center. We observe large differences in performance between newer and older models. Specifically, we find the difference in performance between newer and older models to be significant, positive, and large, and to be roughly similar in magnitude to that estimated for MMLU in Section 2.2. That is, reformulating the test task as multiple choice question answering leads to qualitatively and quantitatively similar confounding to that observed for MMLU. Therefore, newer models overperform on MMLU likely not because of memorization of specific testing data (i.e., due to data contamination or leakage), but rather due to an improved ability for multiple-choice question answering.
Lastly, we adjust for test task training by fine-tuning all models on the MMLU auxiliary training set, and plot their ARC Challenge and HellaSwag scores in Figure 4 right. We no longer find evidence of a large nor a significant difference in performance between newer and older models. Therefore, the proposed adjustment is effective in mitigating the bias introduced by evaluating models via multiple-choice question answering. Notably, performance on ARC and HellaSwag equalizes after fine-tuning on the MMLU auxiliary training set. This indicates that the adjustment data need not closely resemble the test set, but rather the test task.
@@
We evaluate MMLU using the “cloze” methodology instead of the usual multiple-choice prompts. We plot the results in Figure 5 center. With cloze evaluations, the difference in performance between newer and older models is both small and not statistically significant. This suggests that the standard MMLU evaluation conflates knowledge-testing with testing a models’ ability to answer multiple choice questions444For instance, smaller models suffer from particularly strong A-bias for multiple-choice questions (Dominguez-Olmedo et al., 2023).. Newer models therefore attain higher MMLU scores than older models largely because they are better at multiple-choice question answering, and not because they necessarily “know more”.
@@ 4 Implications for model comparisons
So far, we have shown how training on the test task distorts benchmark evaluations. Next, we examine its impact on the relative comparison of model families (Section 4.1), as well as its implications for accurately measuring progress in model capabilities over time (Section 4.2).
@@ 4.1 Comparing model families
We compare the performance of the Pythia, Llama 2, and Qwen 1.5 model families, which likely train on the test task to very different extents. Pythia was trained on the Pile (Gao et al., 2020), a collection of curated datasets that are unlikely to contain much test task data. Llama 2 was trained mostly on web data, which is reasonable to assume may contain more test task data. Lastly, Qwen 1.5 explicitly includes instruction data in its pretraining mixture, thus likely training on the test task to a large extent.
We plot the MMLU and GSM8K scores of the three model families in Figure 6, as well as their adjusted scores (i.e., after fine-tuning on task relevant data). Without adjustment, Qwen 1.5 appears to be the superior model family: it Pareto dominates both the Llama 2 and Pythia models. In contrast, all Pythia models perform no better than random chance, making it unclear whether scaling Pythia offers any benefit at all. After adjustment, however, all three model families exhibit remarkably similar scaling trends. Therefore, after correcting for the confounding of test task training, none of the model families appears superior to the others.
Training on the test task therefore profoundly confounds relative model comparisons. Base models are rarely used “as is” and are generally adapted before deployment. Because of the confounding of training on the test task, performance before adaptation may not reliably predict performance after adaptation.
Interestingly, recent work equates pretraining data quality with downstream benchmark performance (Penedo et al., 2024; Li et al., 2024). Our findings suggest that a major contributing factor to the superior performance of “higher quality” pretraining datasets may be that they simply contain more test task data.
@@ 4.2 Progress in model capabilities
Training on the test task substantially overestimates the progress in benchmark performance per unit of compute achieved by recent model families. In Figure 7 we plot the Pareto frontier of benchmark accuracy against pretraining compute, both for models trained before November 2023 and for all models. We measure progress by considering the area of improvement of the Pareto frontier since November 2023, shaded in green. Without adjustment, the difference between the two Pareto frontiers is large, indicating very substantial progress since November 2023. After adjustment, however, the area of improvement reduces by a sixfold, showing only modest improvements. Therefore, training on the test task strongly overestimates the progress in benchmark performance per unit of compute achieved by recent model families.
On the other hand, recent models tend to be trained on more data than Chinchilla compute-optimal (Hoffmann et al., 2022). Given the Chinchilla scaling laws, it is noteworthy that newer, smaller “over-trained” models match the performance of older, larger ones for the same amount of pretraining compute.
@@ 5 Implications for emergence
Throughout our evaluations, we observe emergent behaviour for MMLU and GSM8K: models perform at near random chance up to a certain scale of pretraining compute, followed by relatively sharper improvements in performance at larger scales (Wei et al., 2022b). After training on the test task, however, emergence for MMLU and GSM8K appears to occur at substantially lower scales. We dedicate this section to more closely investigate the relationship between training on the test task and emergence.
@@
We consider only models trained before November 2023, as we have established that these models train on the test task less than newer models. We evaluate the models at intermediate checkpoints as we fine-tune them on the datasets of task relevant data introduced in Section 2.1. We fit and in Equation 1 to the different intermediate checkpoints, and report in Figure 8 top the corresponding points of emergence . We find that emergence arises at increasingly lower compute regimes as models train on the test task. For MMLU, models exhibit emergence at around FLOPs, the scale of Pythia 6.9B. After training on 64,000 examples, emergence arises around around FLOPs, the scale of Pythia 410M. We observe similar results for GSM8K, see Figure 18 in Appendix E.
@@
The log-linear relationship between pretraining loss and compute is well-established (Kaplan et al., 2020). We observe that, for the compute ranges that we consider, training on the test task increasingly recovers log-linear scaling between pretraining compute and benchmark accuracy. Similarly to the earlier section, we evaluate intermediate checkpoints but instead fit log-linear functions in Figure 8 bottom. We observe that the R2 of the fit improves substantially as the models train on more task-relevant data, jumping from to after training on 64,000 examples. Therefore, after training on the test task almost all of the variation in benchmark accuracy is explained by log-linear scaling of pre-training compute. We observe similar results for GSM8K, see Figure 18 in Appendix E.
Recommendations.
Schaeffer et al. (2024a) argue that emergence appears due to the choice of metric. To mitigate emergence, they suggest to consider Brier score instead of accuracy. We observe, however, that emergence for MMLU does not disappear when using the Brier score, see Figure 5 right, nor that of ARC and HellaSwag when framed as multiple-choice questions, see Figure 17 in Appendix E. We discuss two practical solutions to obtain predictive scaling while maintaining accuracy as the evaluation metric.
For MMLU and multiple-choice benchmarks more broadly, cloze evaluationsc consistently yield smoother and more predictable scaling even when using accuracy as the evaluation metric. Since the purpose of these benchmarks is knowledge-testing more so than testing multiple-choice answering ability, cloze evaluations are preferable insofar predictive scaling at lower compute scales is an important consideration. This recommendation aligns with the concurrent work by Gu et al. (2024).
More broadly, if sufficient task relevant data is available, then training on the test task can result in much more predictable scaling by shifting emergence to smaller compute scales. That is, scaling laws where models across scales are fine-tuned on the same, sufficient task-relevant data prior to evaluation. Such scaling laws correspond to those of specialist models, which for some tasks –e.g., legal annotation (Dominguez-Olmedo et al., 2024)– or purposes –e.g., safety– might be preferable to the scaling laws of generalist models.
@@ 6 Related work
Data contamination or test-set contamination refers to any overlap between the training and the test data such that test results overestimate a model’s generalization performance. The scale and often little curation of present-day pretraining corpora exacerbates data contamination concerns in language model evaluations (Jiang et al., 2024). Consequently, data contamination is usually discussed in the technical reports accompanying model releases (Radford et al., 2019; Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b). However, detecting and preventing data contamination is currently an open problem (Gunasekar et al., 2023; Yang et al., 2023b; Golchin and Surdeanu, 2023). Roberts et al. (2023) and Li and Flanigan (2024) find that models often perform better on datasets that were publicly available during model training. While almost all models that we consider were released after MMLU and GSM8K, we nonetheless find that, controlling for compute, more recent models perform better. These performance gains are unlikely to be driven solely by test set leakage and require additional explanation. In Section 3.2, we find evidence that that training on the test task may be a more dominant factor in benchmark performance than data contamination. A key insight of our work is that models can train on the test task (e.g., multiple-choice question answering) and do much better at a given benchmark (e.g., MMLU) without necessarily seeing any benchmark data. This defies traditionally notions of data contamination, which strictly refer to the leakage of benchmark data (Magar and Schwartz, 2022; Sainz et al., 2023; Dong et al., 2024). Moreover, whereas the core concern with data leakage is that benchmarks may inadvertently find their way to the training data, training on the test task is often a deliberate design choice (e.g., pre-training on instruction data).
Adaptation prior to evaluation.
In the 2010s, language models had to be adapted to different benchmarks using supervised task data (Collobert et al., 2011; Dai and Le, 2015; Devlin et al., 2019). The purpose of such fine-tuned benchmark models was solely to facilitate the relative comparison of base models. Importantly, all models were adapted using the same supervised data (Collobert et al., 2011). With GPT-3 (Brown et al., 2020), few-shot prompting emerged as the dominant paradigm for adapting models to a particular task prior to evaluation (Liang et al., 2023), arguably due to its simplicity relative to fine-tuning.
Bommasani et al. (2021) argue that benchmark evaluations should account for the adaptation resources used by each model (e.g., the adaptation data). Similarly, Liang et al. (2023) argue that the strategy for adapting the models to the benchmark evaluation should be controlled for. Clearly, few-shot prompting is a weaker form of adaptation than training on hundreds of thousands of task examples, as newer models often do. Our proposal of fine-tuning on the same, sufficient amount of task-specific data prior to evaluation aims to effectively control for model adaptation, ensuring that all models are given equal adaptation resources.
@@
The effectiveness of fine-tuning on the training set accompanying LLM benchmarks is well-known (Wei et al., 2022a; Wang et al., 2022; Chung et al., 2024). Consequently, many influential instruction-tuning datasets contain or are partly derived from benchmark train data (Wei et al., 2022a; Honovich et al., 2022; Mukherjee et al., 2023). Li and Flanigan (2024) identify small amounts of benchmark-specific data in the publicly available Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) instruction-tuning sets. Zhou et al. (2023b) empirically analyze the effects of fine-tuning on benchmark-specific data and warn about its impacts on benchmark validity. To circumvent these issues, recent work has focused on indirect indicators of broader data contamination, such as a lack of robustness to task transformations (Wu et al., 2023), or underperformance on benchmarks with novel task combinations (Yu et al., 2023a). In contrast, we find evidence for training on the test task without the need for explicitly identifying specific data points used at training time, or modifying tasks. In addition, our proposed method of fine-tuning on task data prior to evaluation allows us to quantify and correct for the effects of training on the test task on benchmark performance.
@@
Emergent capabilities (Wei et al., 2022b; Ganguli et al., 2022) refer to levels of model performance at large scales that cannot be easily predicted by extrapolating from smaller scales. Wei et al. (2022b) report emergent capabilities for various benchmarks including MMLU and GSM8K (Srivastava et al., 2022). However, Srivastava et al. (2022); Schaeffer et al. (2024b) find that the log-probability of the correct answer often improves smoothly, even when other metrics seem to show emergence. Lu et al. (2023) argue that most emergent capabilities can be explained by in-context-learning. Schaeffer et al. (2024a) argue that emergent capabilities are mostly an artifact of non-linear and discontinuous evaluation metrics like accuracy. In contrast, we find signs of emergence on MMLU even when using continuous metrics like the Brier score. Similarly to our findings, Snell et al. (2024) show that increasingly fine-tuning on the test task shifts the point of emergence to smaller compute scales.